# Regularity Properties of Optimization-Based Controllers

### v4

@@ 5 Forward Invariance Properties of Optimization-Based Controllers
To show invariance invoking Theorem 5.1, one needs to additionally ensure that the closed-loop dynamics has unique solutions. The conditions discussed in Section 4 and summarized in Table 1 can be translated into easily checkable conditions on the objective function, the matrix $A(x)$, and the vector $b(x)$. The following result uses [55, Theorem 3.6] to ensure uniqueness, and therefore forward invariance.
@@
(Sufficient conditions for forward invariance with respect to closed-loop dynamics): Consider the dynamics (27) and the optimization problem (1) where $f\in\mathcal{C}^{1,2}(^{n}\times^{m})$ is strongly convex, and $g(x,u)=b(x)-A(x)u$. Assume
@@
    For all $x\in\mathcal{C}$, there exists $u\in^{m}$ such that $A(x)u>b(x)$, and
@@
    For all $x\in\mathcal{C}$, there is an open set $U_{x}\subset^{n}$ containing $x$ such that, for all $J\subset\{1,\dots,p\}$, the matrix $A_{J}(y)$ has constant rank for all $y\in U_{x}$.
@@
(Minimal Barrier Functions, [72, Theorem 1]): Let $\tilde{F}:^{n}\to^{n}$ be continuous and consider the system $\dot{x}=\tilde{F}(x)$. Let $h:^{n}\to$ be a continuously differentiable function and let $\mathcal{C}=\left\{x\in^{n}\;\mid\;h(x)\geq 0\right\}$ be a nonempty set. If $h$ is a minimal barrier function, cf. [72, Definition 2], then any solution of $\dot{x}=\tilde{F}(x)$ with initial condition in $\mathcal{C}$ remains in $\mathcal{C}$ for all positive times.
@@
Finally, we also note that if $u^{*}$ is discontinuous, the closed-loop system might not have unique solutions and hence the assumptions of Theorem 5.1 will not hold. Therefore, this result cannot be used to guarantee forward invariance of sets. However, the following result gives a sufficient condition for forward invariance of sets under Filippov solutions. It follows as an adaptation of [73, Theorem 1], which gives a sufficient condition for forward invariance of sets under hybrid systems.
@@ 6 Conclusions and Outlook
Regularity properties: Under appropriate constraint qualifications and conditions on the data that defines the optimization problem, we have shown that optimization-based controllers are locally Lipschitz, continuously differentiable, and even analytic. We have also characterized the properties enjoyed by parametric optimizers arising from optimization problems defined by second-order continuously differentiable objective function and constraints, strongly convex objective function, and feasible set with nonempty interior (the same properties as in Robinson’s counterexample). We have shown that, even though such parametric optimizers might not be locally Lipschitz, they enjoy other important regularity properties, like point-Lipschitz continuity. Even if the optimization-based controller is discontinuous, under appropriate conditions on the optimization problem data, we have shown that it is measurable and locally bounded.
@@
Outlook: The results presented in this work show that the regularity properties of optimization-based controllers are determined by the smoothness/convexity and constraint qualification properties of the optimization problems defining them. This opens the door to the possibility of designing optimization problems with the appropriate conditions and constraint qualification properties in order to endow the associated optimization-based controller with certain desired regularity properties. For example, in the context of safety-critical control, it is sufficient to find safe sets and control barrier functions for those sets to guarantee that an associated control barrier function based controller is locally Lipschitz. In certain cases, one can guarantee that these control objectives are obtained without continuity or even uniqueness of solutions to the resulting closed-loop systems. Characterizing conditions on the objective function and constraints to ensure that control objectives are achieved even in the absence of local Lipschitz continuity is an important direction for future work. In particular, understanding one-sided Lipschitzness in the context of parametric optimization is important in the context of safety-critical control, since uniqueness of solutions allows one to verify forward-invariance of a safe set via Nagumo’s Theorem. Besides the examples mentioned in Section 1, the relevance of the results presented herein also applies to other areas of systems and control, where controllers need to be designed with desirable regularity properties, such as backstepping [74], where virtual controllers need to be differentiated at the intermediate layers to construct a composite Lyapunov or barrier [50] function for the composite system. Other ideas for future work include improving further the results presented in this paper to more specific classes of problems, such as CBF-based quadratic programs or second-order convex programs, as well as MPC-based controllers. We also plan to relax the differentiability and convexity assumptions considered throughout this paper and give regularity results for set-valued optimizers using the theory of continuous selections.
@@ Acknowledgments
This work was partially supported by ARL-W911NF-22-2-0231. The authors would like to thank E. Dall’Anese for multiple conversations on optimization-based controllers, S. Ganguly for pointing out typos in an earlier version of the manuscript, an anonymous reviewer of the paper [44] for various comments regarding Robinson’s counterexample, and M. Alyaseen for a discussion regarding Example 3.6.

### v5

@@ 5 Forward Invariance Properties of Optimization-Based Controllers
To show invariance invoking Theorem 5.1, one needs to additionally ensure that the closed-loop dynamics has unique solutions. The conditions discussed in Section 4 and summarized in Table 1 can be translated into easily checkable conditions on the objective function, the matrix $A(x)$, and the vector $b(x)$. The following result uses [54, Theorem 3.6] to ensure uniqueness, and therefore forward invariance.
@@
(Sufficient conditions for forward invariance with respect to closed-loop dynamics): Consider the dynamics (27) and the optimization problem (1) where $f\in\mathcal{C}^{1,2}(^{n}\times^{m})$ is strictly convex, and $g(x,u)=b(x)-A(x)u$. Assume
@@
    For all $x\in\mathcal{C}$, there exists $u\in^{m}$ such that $A(x)u>b(x)$, and (1) has at least one minimizer.
@@
    For all $x\in\mathcal{C}$, there is an open set $U_{x}\subset^{n}$ containing $x$ such that, for any subset $J$ in $\{1,\dots,p\}$, the matrix $A_{J}(y)$ has constant rank for all $y\in U_{x}$.
@@
(Minimal Barrier Functions, [74, Theorem 1]): Let $\tilde{F}:^{n}\to^{n}$ be continuous and consider the system $\dot{x}=\tilde{F}(x)$. Let $h:^{n}\to$ be a continuously differentiable function and let $\mathcal{C}=\left\{x\in^{n}\;\mid\;h(x)\geq 0\right\}$ be a nonempty set. If $h$ is a minimal barrier function, cf. [74, Definition 2], then any solution of $\dot{x}=\tilde{F}(x)$ with initial condition in $\mathcal{C}$ remains in $\mathcal{C}$ for all positive times.
@@
Finally, we also note that if $u^{*}$ is discontinuous, the closed-loop system might not have unique solutions and hence the assumptions of Theorem 5.1 will not hold. Therefore, this result cannot be used to guarantee forward invariance of sets. However, the following result gives a sufficient condition for forward invariance of sets under Filippov solutions. It follows as an adaptation of [75, Theorem 1], which gives a sufficient condition for forward invariance of sets under hybrid systems.
@@ 6 Conclusions and Outlook
Regularity properties: Under appropriate constraint qualifications and conditions on the data that defines the optimization problem, we have shown that optimization-based controllers are locally Lipschitz, continuously differentiable, and even analytic. We have also characterized the properties enjoyed by parametric optimizers arising from optimization problems defined by second-order continuously differentiable objective function and constraints, strictly (or strongly) convex objective function, and feasible set with nonempty interior (the same properties as in Robinson’s counterexample). We have shown that, even though such parametric optimizers might not be locally Lipschitz, they enjoy other important regularity properties, like point-Lipschitz continuity. Even if the optimization-based controller is discontinuous, under appropriate conditions on the optimization problem data, we have shown that it is measurable and locally bounded.
@@
Outlook: The results presented in this work show that the regularity properties of optimization-based controllers are determined by the smoothness/convexity and constraint qualification properties of the optimization problems defining them. This opens the door to the possibility of designing optimization problems with the appropriate conditions and constraint qualification properties in order to endow the associated optimization-based controller with certain desired regularity properties. For example, in the context of safety-critical control, it is sufficient to find safe sets and control barrier functions for those sets to guarantee that an associated control barrier function based controller is locally Lipschitz. In certain cases, one can guarantee that these control objectives are obtained without continuity or even uniqueness of solutions to the resulting closed-loop systems. Characterizing conditions on the objective function and constraints to ensure that control objectives are achieved even in the absence of local Lipschitz continuity is an important direction for future work. In particular, understanding one-sided Lipschitzness in the context of parametric optimization is important in the context of safety-critical control, since uniqueness of solutions allows one to verify forward-invariance of a safe set via Nagumo’s Theorem. Besides the examples mentioned in Section 1, the relevance of the results presented herein also applies to other areas of systems and control, where controllers need to be designed with desirable regularity properties, such as backstepping [76], where virtual controllers need to be differentiated at the intermediate layers to construct a composite Lyapunov or barrier [49] function for the composite system. Other ideas for future work include improving further the results presented in this paper to more specific classes of problems, such as CBF-based quadratic programs or second-order convex programs, as well as MPC-based controllers. We also plan to relax the differentiability and convexity assumptions considered throughout this paper and give regularity results for set-valued optimizers using the theory of continuous selections [77].
@@ Acknowledgments
This work was partially supported by ARL-W911NF-22-2-0231. The authors would like to thank E. Dall’Anese and Y. Chen for multiple conversations on optimization-based controllers, S. Ganguly for pointing out typos in an earlier version of the manuscript, an anonymous reviewer of the paper [43] for various comments regarding Robinson’s counterexample, and M. Alyaseen for a discussion regarding Example 3.6.